OpenStudy (anonymous):

The perimeter of a rectangle is 70 cm. The width is (4x - 1)cm and the length is (6x - 4)cm. What is the length of the rectangle? The perimeter of a rectangle is 70 cm. The width is (4x - 1)cm and the length is (6x - 4)cm. What is the length of the rectangle? @Mathematics

OpenStudy (anonymous):

width = 15 length = 20

hero (hero):

Sheg, how come you never post steps?

OpenStudy (cathyangs):

so then 2(4x-1)+2(6x-4)=70 this simplifies to: 8x-2+12x-8=70 then 20x=80 x=4. 24-4=20 cm length.

OpenStudy (anonymous):

oh so sorry dear hero will post it

OpenStudy (anonymous):

hero if the asker asks me the steps then only i post the steps

OpenStudy (anonymous):

2(Length + Width) = Perimeter 2{ (6x-4) + (4x-1)} = 70 10x - 5 = 35 10x = 40 x = 4 Width = 4*4 - 1 = 15 Length = 6*4 - 4 = 20

OpenStudy (anonymous):

hero satisfied

hero (hero):

You should do that more often

OpenStudy (anonymous):

ok bro as u wish